OpenStudy (anonymous):

In which of these pairs of numbers is 1.75 greater than the first number but smaller than the second number? 0 and 1 1 and 3/2 3/2 and 1 9/10 1 9/10 and 2

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

i dont know i cant do fractions at all.

OpenStudy (anonymous):

0 and 1=0.00 and 1.00 1 and 3/2=1.00 and 1.50 3/2 and 1 9/10=1.50 and 1.90 1 9/10 and 2=1.90 and 2.00

OpenStudy (anonymous):

then just put in the question

OpenStudy (anonymous):

ok got it